OpenStudy (anonymous):

Create your own example of a mathematical phrase you might hear in the real world and provide the translation for it. Be sure to identify what part of the phrase is represented by your variable.

OpenStudy (anonymous):

the sum of the cash in our pockets is five dollars and ninety five cents

OpenStudy (anonymous):

C= coins in my pocket and B= bills in your pocket therefore C+B=5.95$

OpenStudy (anonymous):

thanks!

OpenStudy (jamesj):

"And there's 50 cents change." Let P = price in $, M = money paid in $. Then M - P = 1/2